Building A is 190 feet shorter than Building B. The total height of the two buildings is 1520 feet. Find the height of each building.
What is the height of Building A?
What is the height of Building B?

Answer by fractalier(6550)

From the facts we have
A = B - 190
and
A + B = 1520
Couple ways to solve this, but let's substitute the first into the second...so
B - 190 + B = 1520
2B - 190 = 1520
2B = 1710
B = 855 feet
so that A is 190 shorter or
A = 665 feet